# Large Language Model Agent for Modular Task Execution in Drug Discovery

### v2

@@ Abstract
We present a modular framework powered by large language models (LLMs) that automates and streamlines key tasks across the early-stage computational drug discovery pipeline. By combining LLM reasoning with domain-specific tools, the framework performs biomedical data retrieval, domain-specific question answering, molecular generation, property prediction, property-aware molecular refinement, and 3D protein–ligand structure generation. In a case study targeting BCL-2 in lymphocytic leukemia, the agent autonomously retrieved relevant biomolecular information, including FASTA sequences, SMILES representations, and literature, and answered mechanistic questions with improved contextual accuracy compared to standard LLMs. It then generated chemically diverse seed molecules and predicted 67 ADMET-related properties, which guided iterative molecular refinement. Across two refinement rounds, the number of molecules with QED >> 0.6 increased from 34 to 55. The number of molecules satisfying empirical drug-likeness filters also rose; for example, compliance with the Ghose filter increased from 32 to 55 within a pool of 100 molecules. The framework also employed Boltz-2 to generate 3D protein–ligand complexes and provide rapid binding affinity estimates for candidate compounds. These results demonstrate that the approach effectively supports molecular screening, prioritization, and structure evaluation. Its modular design enables flexible integration of evolving tools and models, providing a scalable foundation for AI-assisted therapeutic discovery.
@@ 1 Introduction
The discovery of drug-like molecules for treating specific diseases is fundamental to therapeutic innovation and the advancement of public health. However, identifying viable candidates is a highly complex and resource-intensive process. It requires navigating an enormous chemical space, integrating heterogeneous biological and chemical data sources, and conducting iterative experimental validation. Empirically, drug development spans 10-15 years and can cost \$1.5 - 2.5 billion per approved therapy. Much of this burden arises from attrition in Phase II and III clinical trials, where insufficient efficacy or safety issues often lead to late-stage failures 1, 2, 3. A key strategy to reduce this attrition is to identify high-quality candidates as early as possible, thereby improving the likelihood of clinical success. At the same time, the chemical search space is astronomically large, with estimates of more than $10^{60}$ drug-like molecules, making it impossible to explore effectively with traditional trial-and-error approaches 4, 5.
@@
To mitigate these challenges, computational approaches have become essential in early-stage drug development. Methods such as molecular docking 6, 7, 8, quantitative structure–activity relationship (QSAR) modeling 9, 10, 11, and molecular dynamics simulations 12, 13 have significantly improved the speed and accuracy of compound screening and optimization. More recently, machine learning (ML) techniques have expanded this toolkit by enabling predictive modeling of pharmacokinetic and pharmacodynamic properties, de novo molecular generation, and multitask learning across diverse biomedical tasks14, 15, 16, 17. Notably, AlphaFold has achieved high-accuracy protein structure prediction 18, 19, 20, while Graph Neural Networks (GNNs) and transformer-based models improve property prediction 21, 22, 23, 24. In parallel, generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models (e.g., JT-VAE, MolGAN, MolDiff) enable de novo or target-specific molecular design 25, 26, 27. Collectively, these methods increase the scalability and automation potential of modern drug discovery workflows.
@@
Despite this progress, most computational methods remain task-specific and require manual orchestration by domain experts. However, drug discovery is inherently a multi-step, interdependent process that requires seamless integration across diverse tasks 28, 29, 30. This fragmented implementation limits the scalability and efficiency of current pipelines. As the demand for faster and more cost-effective drug screening grows, there is an urgent need for unified, intelligent platforms capable of autonomously coordinating these tasks while supporting expert decision-making.
@@
Large Language Models (LLMs) offer a compelling solution to this need. Trained on massive corpora of natural language data, LLMs exhibit strong reasoning capabilities and domain-agnostic knowledge. When augmented with external tools, such as domain-specific plugins, APIs, and software libraries, LLM-based agents can overcome the limitations of general-purpose language models and act as interpretable, flexible controllers of scientific workflows 31. LLM agents have recently been applied to automate diverse aspects of scientific discovery, including experimental design, material synthesis planning, and data analysis 32, 33, 34, 35, 36. For instance, Boiko et al. introduced Coscientist, an LLM agent capable of autonomously planning and executing chemistry experiments, significantly enhancing lab productivity while reducing human intervention 37. Similarly, the dZiner framework leverages LLM agents for molecular design through iterative reasoning and structure-based optimization38.
@@
Furthermore, CACTUS39 demonstrated the use of an LLM to wrap cheminformatics utilities for evaluating properties such as Mwt, LogP, TPSA, and drug-likeness filters, providing a practical but narrowly scoped toolset. In contrast, SciToolAgent40 introduced a knowledge-graph–driven orchestration layer that allows an LLM to select from hundreds of scientific tools across domains ranging from biological to materials, illustrating scalability but lacking a dedicated workflow for therapeutic design. Building on these directions, DrugPilot41 emphasizes parameterized reasoning for multi-stage discovery tasks, though its scope remains tied to a fixed set of model zoo tasks and benchmark datasets, without iterative molecular refinement, limiting its ability to improve candidate quality over successive rounds.
@@
In this work, we introduce AgentD, an LLM-powered agent framework designed to support and streamline the drug discovery pipeline. The agent performs a wide range of essential tasks, including biomedical data retrieval from structured databases and unstructured web sources, answering domain-specific scientific queries, generating seed molecule libraries via SMILES-based generative models, predicting a broad spectrum of drug-relevant properties, refining molecular representations to improve drug-likeness, and generating 3D molecular structures for downstream analysis. Our results demonstrate that this agent-driven framework streamlines early-phase drug discovery and serves as a flexible foundation for scalable, AI-assisted therapeutic development. Furthermore, its modular architecture allows for continual improvement as more advanced tools and models become available.
2 Agent Design
@@ 2.1 Task Modules
AgentD performs six essential tasks across the drug discovery pipeline, as illustrated in Figure 1. The agent leverages large language models from providers such as OpenAI and Anthropic; in this study, we primarily use OpenAI’s GPT-4o as both the central reasoning engine and the natural language interface. By integrating domain-specific tools and databases, AgentD coordinates a wide range of activities - from data retrieval and molecular generation to property evaluation and structure prediction. The primary tool components supporting each task module are summarized in Table 1.
Figure 1: Overview of task modules supported by AgentD. In the question answering and molecule refinement tasks, outputs are generated directly by the language model, whereas in other tasks, the final results are produced by integrated external tools.
Data Extraction. AgentD is capable of retrieving biomedical information from both structured and unstructured sources. Given a query - such as identifying known drugs associated with a specific protein–disease context - the agent retrieves the protein’s FASTA sequence from the UniProt database 42, searches for relevant drug names via the web, and extracts corresponding SMILES representations from the ChEMBL database 43. Additionally, AgentD autonomously constructs keyword-based queries to download relevant open-access scientific literature from Semantic Scholar, providing a broader context for downstream reasoning and decision-making.
@@
SMILES Generation. Constructing a diverse and chemically relevant seed molecule pool is essential for effective early-stage virtual screening. AgentD generates seed molecules in SMILES format using external generative models. In this study, we incorporate two such models: REINVENT 44, 45, which supports de novo molecule generation without requiring input SMILES, and Mol2Mol 46, which performs conditional generation to produce molecules structurally similar to a given input SMILES. This dual capability enables both exploration and exploitation in the molecular search space.
Property Prediction. For each candidate molecule, AgentD predicts key pharmacologically relevant properties, including ADMET (absorption, distribution, metabolism, excretion, and toxicity) profiles and binding affinity (e.g., $pK_{d}$). ADMET prediction is performed using the Deep-PK API 47, which accepts SMILES strings as input. For binding affinity estimation, the BAPULM model 48 is used, which operates on both the SMILES and the protein’s FASTA sequence. These predictions help prioritize compounds based on both efficacy and safety.
Molecule Refinement. Based on the predicted properties, AgentD can identify molecular shortcomings such as toxicity or poor permeability, and propose targeted structural modifications to improve attributes like solubility and metabolic stability. This SMILES refinement is carried out solely through the LLM’s internal reasoning and built-in chemical knowledge, without additional model-based tools 38.
Structure Generation. AgentD can generate 3D structures of protein–ligand complexes using Boltz-2 as an external tool 49, 50. This process produces candidate complex structures along with associated binding metrics such as IC50 values and inhibitor probability. These structures can serve as inputs for downstream computational tasks such as docking simulations or molecular dynamics, offering deeper insight into the biophysical interactions of the drug candidates.
@@
Data Extraction UniProt, ChEMBL, Serper, Semantic Scholar APIs
@@
SMILES Generation REINVENT, Mol2Mol
Property Prediction DeepPK API, BAPULM
@@
Structure Generation Boltz-2
*Adapted from the dZiner framework 38.
@@ 2.2 Workflow
All six task modules in AgentD are designed to support key components of the drug discovery pipeline, as illustrated in Figure 2. In this study, we demonstrate the workflow using a case study involving BCL-2, a well-characterized therapeutic target in lymphocytic leukemia. The process begins with a user-provided query - for instance, identifying drug molecules that target a specific disease-related protein. Through the data extraction module, AgentD retrieves the protein’s FASTA sequence from UniProt and identifies known drugs using sources such as ChEMBL and Google search APIs.
The SMILES of a known drug serves as a starting point for the SMILES generation task, which builds a chemically diverse library of candidate molecules using generative models. These molecules are then evaluated using property prediction tools to estimate key pharmacological characteristics such as ADMET profiles and binding affinity. Based on these predicted properties, AgentD proposes SMILES-level modifications to improve attributes like solubility, toxicity, and metabolic stability. These refinements rely on the language model’s internal reasoning rather than external optimization tools and help enrich the candidate pool.
Following refinement, domain-specific criteria can be applied by the user to select promising compounds. For these, the structure generation task creates 3D protein–ligand complexes using the ligand SMILES and target protein sequence. This enables more detailed downstream analyses. The structure generation step also outputs auxiliary metrics like predicted IC50 and inhibitor probability as proxies for binding strength. While these estimates, derived from Boltz-2, offer useful guidance, they should be interpreted cautiously, as they are not highly accurate.
@@
Figure 2: Workflow illustrating how each task module within AgentD supports the drug discovery pipeline.
@@ 3.1 Data Extraction
Reliable access to protein sequences and ligand structures is essential for computational tasks such as virtual screening, molecular modeling, and property prediction. To support this capability, AgentD integrates web search functionality and database API access as tool modules. When provided with a user query specifying a target protein and associated disease, the agent is instructed to: (i) retrieve the protein’s FASTA sequence, (ii) identify existing drugs relevant to the specified target–disease context, and (iii) download related open-access scientific literature.
The agent queries the UniProt database for the protein sequence and uses web search tools (via the Serper API) to identify relevant drug molecules. Once a drug is identified, its SMILES representation is retrieved from the ChEMBL database. In parallel, the agent generates keyword-based queries to download supporting publications from Semantic Scholar. To demonstrate this functionality, the agent was tested on example queries including BCL-2 for lymphocytic leukemia, JAK2 for myelofibrosis, and KIT for renal cell carcinoma. The results, summarized in Figure 3, show that the agent successfully retrieved key molecular data, such as identifying venetoclax as a BCL-2 inhibitor.
@@
Figure 3: Examples of the data extraction. Given a user query specifying a target protein and associated disease, AgentD retrieves the protein’s FASTA sequence and identifies existing drug molecules related to the query.
@@ 3.2 Domain-Specific Question Answering
Figure 4: Comparison of question answering performance between AgentD with RAG and the standard GPT-4o model. Green highlights indicate grounded, context-relevant information supported by the reference paper, while grey highlights denote generic content not directly related to the question.
We evaluated this capability using a representative study by Weller et al., which describes how venetoclax activates the integrated stress response (ISR), leading to NOXA upregulation and MCL-1 inhibition 51. Based on this paper, we formulated domain-specific questions and compared the responses generated by AgentD with RAG to those from the standard GPT-4o model (Figure 4).
AgentD consistently generates more detailed and mechanistically accurate responses. For example, when asked about the synergy mechanism between venetoclax and proteasome inhibitors, the RAG-augmented answer includes key elements such as ISR activation, ATF3/ATF4-mediated transcription of NOXA, and downstream effects involving BIM, BAX/BAK activation, and mitochondrial membrane permeabilization,closely aligning with the mechanistic explanation presented in the source paper (Figure 4a). In contrast, GPT-4o provides a more generic explanation, omitting critical components such as the ISR pathway and transcriptional regulators.
@@
Table 2: Comparison of BERTScores between AgentD (with RAG) and standard GPT-4o (without RAG). Values are averaged across six question–answer pairs. Results for two representative pairs are shown in Figure 4, and additional examples are provided in Supplementary Information S1.
@@
AgentD (w/ RAG) 0.720 $\pm$ 0.048 0.687 $\pm$ 0.057 0.702 $\pm$ 0.041
Standard GPT4o (w/o RAG) 0.602 $\pm$ 0.023 0.669 $\pm$ 0.033 0.633 $\pm$ 0.022
To quantitatively assess these observed differences, we measured the semantic similarity between responses generated with and without RAG and human-annotated reference answers derived from the literature using BERTScore 52. The reference answers were curated by the authors through careful reading of the relevant source papers 51, 53. BERTScore quantifies contextual similarity by comparing the embeddings of model-generated and reference texts. We employed SciBERT 54 as the embedding encoder because it is pre-trained on a large corpus of scientific publications from Semantic Scholar, capturing richer representations of biomedical and technical terminology than general-domain encoders such as RoBERTa 41. This setup provides a biologically meaningful and context-aware framework for evaluating the factual basis of model outputs in literature-based question-answering.
As shown in Table 2, AgentD with RAG achieved notably higher semantic fidelity to the reference literature (F1 = 0.70 $\pm$ 0.04) compared to standard GPT-4o (F1 = 0.63 $\pm$ 0.02), with the largest improvement observed in Precision—indicating that RAG effectively filters unsupported or overly generic content. Specifically, for the question shown in Figure 4a, the BERTScores obtained with RAG were Precision = 0.675, Recall = 0.697, and F1 = 0.686, while the corresponding scores without RAG were 0.616, 0.666, and 0.640, respectively. Similarly, for the question in Figure 4b, the BERTScores with RAG were Precision = 0.616, Recall = 0.666, and F1 = 0.640, compared to 0.623, 0.694, and 0.660 without RAG. Consistent with other test cases, these examples demonstrate that incorporating RAG leads to improved semantic alignment and factual grounding.

### v3

@@ Abstract
We present a modular framework powered by large language models (LLMs) that automates and streamlines key tasks across the early-stage computational drug discovery pipeline. By combining LLM reasoning with domain-specific tools, the framework performs biomedical data retrieval, literature-grounded question answering via retrieval-augmented generation, molecular generation, multi-property prediction, property-aware molecular refinement, and 3D protein–ligand structure generation. The agent autonomously retrieved relevant biomolecular information, including FASTA sequences, SMILES representations, and literature, and answered mechanistic questions with improved contextual accuracy compared to standard LLMs. It then generated chemically diverse seed molecules and predicted 75 properties, including ADMET-related and general physicochemical descriptors, which guided iterative molecular refinement. Across two refinement rounds, the number of molecules with QED >> 0.6 increased from 34 to 55. The number of molecules satisfying empirical drug-likeness filters also rose; for example, compliance with the Ghose filter increased from 32 to 55 within a pool of 100 molecules. The framework also employed Boltz-2 to generate 3D protein–ligand complexes and provide rapid binding affinity estimates for candidate compounds. These results demonstrate that the approach effectively supports molecular screening, prioritization, and structure evaluation. Its modular design enables flexible integration of evolving tools and models, providing a scalable foundation for AI-assisted therapeutic discovery.
@@ 1 Introduction
The discovery of drug-like molecules is fundamental to therapeutic innovation and the advancement of public health. However, identifying viable candidates is a highly complex and resource-intensive process. It requires navigating an enormous chemical space, integrating heterogeneous biological and chemical data sources, and conducting iterative experimental validation. Empirically, drug development spans 10-15 years and can cost \$1.5 - 2.5 billion per approved therapy. Much of this burden arises from attrition in Phase II and III clinical trials, where insufficient efficacy or safety issues often lead to late-stage failures Paul2010-sr, Roland2024-pw, Subbiah2023-tw. A key strategy to reduce this attrition is to identify high-quality candidates as early as possible, thereby improving the likelihood of clinical success. At the same time, the chemical search space is astronomically large, with estimates of more than $10^{60}$ drug-like molecules, making it impossible to explore effectively with traditional trial-and-error approaches Bohacek1996, Polishchuk2013.
@@
To mitigate these challenges, computational approaches have become essential in early-stage drug development. Methods such as molecular docking Fan2019, autodock_vina, docking_software, quantitative structure–activity relationship (QSAR) modeling qsar, qsar_2, dl_qsar, and molecular dynamics simulations md_for_protein_1, md_for_protein_2, gradnav have significantly improved the speed and accuracy of compound screening and optimization. More recently, machine learning (ML) techniques have expanded this toolkit by enabling predictive modeling of pharmacokinetic and pharmacodynamic properties, de novo molecular generation, and multitask learning across diverse biomedical tasksbadrinarayanan2025mofgptgenerativedesignmetalorganic, peptidebert, idpbert, multipeptide. AlphaFold has achieved high-accuracy protein structure prediction alphafold, alphafold2, alphafold3. In parallel, generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models (e.g., JT-VAE, MolGAN, MolDiff) enable de novo or target-specific molecular design jt_vae, molgan, moldiff. Also, Graph Neural Networks (GNNs) and transformer-based models improve property prediction admetlab2, moleculenet, molprop, deeptox. Collectively, these methods increase the scalability and automation potential of modern drug discovery workflows.
@@
Despite this progress, most computational methods remain task-specific and require manual orchestration by domain experts. Drug discovery is inherently a multi-step, interdependent process that requires seamless integration across diverse tasks Berdigaliyev2020, Hughes2011, SINHA201819. This fragmented implementation limits the scalability and efficiency of current pipelines. As the demand for faster and more cost-effective drug screening grows, there is an urgent need for unified, intelligent platforms capable of autonomously coordinating these tasks while supporting expert decision-making.
@@
Large Language Models (LLMs) offer a compelling solution to this need. Trained on massive corpora of natural language data, LLMs exhibit strong reasoning capabilities and domain-agnostic knowledge. When augmented with external tools, such as domain-specific plugins, APIs, and software libraries, LLM-based agents can overcome the limitations of general-purpose language models and act as interpretable, flexible controllers of scientific workflows aiscientist, coscientist.
@@
LLM agents have recently been applied to automate diverse aspects of scientific discovery, including experimental design, material synthesis planning, and data analysis chemcrow, adsorb_agent, llm_design, llm3dprint, matsci_agent. For instance, the dZiner framework leverages LLM agents for molecular design through iterative reasoning and structure-based optimizationdziner. CACTUSMcNaughton2024-bp demonstrated the use of an LLM to wrap cheminformatics utilities for evaluating properties such as Mwt, LogP, TPSA, and drug-likeness filters, providing a practical but narrowly scoped toolset. In contrast, SciToolAgentDing2025-he introduced a knowledge-graph–driven orchestration layer that allows an LLM to select from hundreds of scientific tools across domains ranging from biological to materials, illustrating scalability but lacking a dedicated workflow for therapeutic design. Building on these directions, DrugPilotLi2025-iy focuses on orchestrating multi-stage discovery workflows using parameterized reasoning, primarily demonstrated on established model-zoo tasks and benchmark datasets rather than end-to-end molecular design. A more detailed overview of the comparative landscape of agentic frameworks is provided in Supplementary Information Section 9.
@@
We introduce AgentD, an LLM-powered agent framework designed to support and streamline the drug discovery pipeline. The agent performs a wide range of important tasks, including biomedical data retrieval from structured databases and unstructured web sources, answering domain-specific scientific queries, generating seed molecule libraries via SMILES-based generative models, predicting a broad spectrum of drug-relevant properties, refining molecular representations to improve drug-likeness, and generating 3D molecular structures for downstream analysis. Our results demonstrate that this agent-driven framework streamlines early-phase drug discovery and serves as a flexible foundation for scalable, AI-assisted therapeutic development. Its modular architecture allows for continual improvement as more advanced tools and models become available.
2 Agentic Workflow
@@ 2.1 Task Modules
AgentD performs six essential tasks across the drug discovery pipeline, as illustrated in Figure 1a. From a code-design perspective, the agent integrates LLMs sourced from OpenAI, Anthropic, and DeepSeek. These models function both as the core reasoning module and as the natural-language interaction layer. In the present implementation, AgentD is primarily developed and executed using OpenAI’s GPT-4o as the main language model. By integrating domain-specific tools and databases, AgentD coordinates a wide range of activities - from data retrieval and molecular generation to property evaluation and structure prediction. The primary tool components supporting each task module are summarized in Table 1.
Figure 1: AgentD overview. a Task modules supported by AgentD. In question-answering and molecular-refinement tasks, outputs are generated directly by the language model, while other tasks rely on external computational tools for final results. b Workflow illustrating how each task module contributes to the overall drug discovery pipeline.
Data Extraction. AgentD is capable of retrieving biomedical information from both structured and unstructured sources. Given a query, such as identifying known drugs associated with a specific protein–disease context, the agent retrieves the protein’s FASTA sequence from UniProt uniprot, searches for relevant drug names on the web, and extracts corresponding SMILES representations from ChEMBL chembl. Additionally, AgentD autonomously constructs keyword-based queries to download relevant open-access scientific literature from Semantic Scholar, providing a broader context for downstream reasoning and decision-making.
@@
SMILES Generation. Constructing a diverse and chemically relevant seed molecule pool is essential for effective early-stage virtual screening. AgentD generates seed molecules in SMILES format using external generative models. In this study, we incorporate two such models: REINVENT reinvent2, reinvent4, which supports de novo molecule generation without requiring input SMILES, and Mol2Mol mol2mol, which performs conditional generation to produce molecules structurally similar to a given input SMILES. This dual capability enables both exploration and exploitation in the molecular search space.
Property Prediction. For each candidate molecule, AgentD predicts key pharmacologically relevant properties, including ADMET (absorption, distribution, metabolism, excretion, and toxicity) profiles and binding affinity (e.g., $pK_{d}$). ADMET prediction is performed using the Deep-PK API deep_pk, which accepts SMILES strings as input. For binding affinity estimation, the BAPULM model bapulm is used, which operates on both the SMILES and the protein’s FASTA sequence. These predictions help prioritize compounds based on both efficacy and safety.
Molecule Refinement. Based on the predicted properties, AgentD can identify molecular shortcomings such as toxicity or poor permeability, and propose targeted structural modifications to improve attributes like solubility and metabolic stability. This SMILES refinement is carried out solely through the LLM’s internal reasoning and built-in chemical knowledge, without additional model-based tools dziner. This task is carried out on top of the property prediction stage.
Structure Generation. AgentD can generate 3D structures of protein–ligand complexes using Boltz-2 as an external tool wohlwend2024boltz1, passaro2025boltz2. This process produces candidate complex structures along with associated binding metrics such as IC50 values and inhibitor probability. These structures can serve as inputs for downstream computational tasks such as docking simulations or molecular dynamics, offering deeper insight into the biophysical interactions of the drug candidates.
@@
Data Extraction UniProt uniprot, ChEMBL chembl, Serper serperAPI, Semantic Scholar APIs
@@
SMILES Generation REINVENT reinvent4, Mol2Mol mol2mol
Property Prediction DeepPK API deep_pk, BAPULM bapulm
@@
Structure Generation Boltz-2 passaro2025boltz2
*Adapted from the dZiner framework dziner.
@@ 2.2 Workflow
All six task modules in AgentD are designed to support key components of the drug discovery pipeline, as illustrated in Figure 1b. The process begins with a user-provided query, such as identifying drug molecules that target a specific disease-related protein. Through the data extraction module, AgentD retrieves the protein’s FASTA sequence from UniProt and identifies known drugs using sources such as ChEMBL and Google search APIs. In this study, we demonstrate the workflow using three representative therapeutic targets: BCL-2 for lymphocytic leukemia, JAK-2 for myelofibrosis, and thrombin for thrombosis. The BCL-2 case is discussed in detail in the main manuscript, while results for the other two targets are provided in Supplementary Information Sections 5 and 6.
The AgentD framework checks whether known drug molecules exist for a given target or disease. If available, these molecules serve as starting points for subsequent molecular exploration. Because AgentD includes a de novo generation module, the workflow can still proceed even when no such molecules are found. The SMILES generation task builds a chemically diverse library of candidate molecules using generative models. These molecules are then evaluated using property prediction tools to estimate key pharmacological characteristics such as ADMET profiles, general physicochemical properties, and binding affinity. Based on these predicted properties, AgentD proposes SMILES-level modifications to improve attributes like solubility, toxicity, and metabolic stability. These refinements rely on the language model’s internal reasoning rather than external optimization tools and help enrich the candidate pool.
Following refinement, domain-specific criteria can be applied by the user to select promising compounds. For these, the structure generation task creates 3D protein–ligand complexes using the ligand SMILES and target protein sequence. This enables more detailed downstream analyses. The structure generation step also outputs auxiliary metrics like predicted IC50 and inhibitor probability as proxies for binding strength. Although these estimates from Boltz-2 provide useful initial guidance, they should ultimately be validated using more reliable computational methods, such as molecular dynamics simulations. Additional uncertainty information, including Boltz-2–derived affinity estimates, is provided in Supplementary Information Section 10.
@@ 3.1 Data Extraction
AgentD integrates web search capabilities and database API access as tool modules for extracting protein and molecule data. When provided with a user query specifying a target protein and associated disease, the agent is instructed to: (i) retrieve the protein’s FASTA sequence, (ii) identify existing drugs relevant to the specified target–disease context, and (iii) download related open-access scientific literature. The agent was tested on example queries including BCL-2 for lymphocytic leukemia, JAK2 for myelofibrosis, and thrombin for thrombosis. The results, summarized in Figure 2, show that the agent successfully retrieved key molecular data, such as identifying venetoclax as a BCL-2 inhibitor.
@@
Figure 2: Examples of data extraction. Given a user query specifying a target protein and associated disease, AgentD retrieves the protein’s FASTA sequence and identifies known drug molecules relevant to the query. a BCL-2 for lymphocytic leukemia, b JAK-2 for myelofibrosis, and c thrombin for thrombosis.
@@ 3.2 Domain-Specific Question Answering
Figure 3: Comparison of question-answering performance between AgentD with RAG and the Vanilla GPT-4o model. Green highlights indicate grounded, context-relevant information supported by the reference paper, while grey highlights denote generic content not directly related to the question. a Mechanism of ABT-199/venetoclax synergy with proteasome inhibitors. b Effect of ABT-199–induced integrated stress response on NOXA expression.
We evaluated this capability using a representative study by Weller et al., which describes how venetoclax activates the integrated stress response (ISR), leading to NOXA upregulation and MCL-1 inhibition qna_ex1. Based on this paper, we formulated domain-specific questions and compared the responses generated by AgentD with RAG to those from the vanilla GPT-4o model (Figure 3).
AgentD consistently generates more detailed and mechanistically accurate responses. For example, when asked about the synergy mechanism between venetoclax and proteasome inhibitors, the RAG-augmented answer includes key elements such as ISR activation, ATF3/ATF4-mediated transcription of NOXA, and downstream effects involving BIM, BAX/BAK activation, and mitochondrial membrane permeabilization,closely aligning with the mechanistic explanation presented in the source paper (Figure 3a). In contrast, vanilla GPT-4o provides a more generic explanation, omitting critical components such as the ISR pathway and transcriptional regulators.
@@
Table 2: Comparison of BERTScores between AgentD (with RAG) and vanilla language models (without RAG). Values are averaged across six question–answer pairs. Results for two representative pairs are shown in Figure 3, and additional examples are provided in Supplementary Information Section 1.
@@
AgentD (GPT-4o, w/ RAG) 0.719 $\pm$ 0.048 0.686 $\pm$ 0.064 0.701 $\pm$ 0.046
Vanilla GPT-4o (w/o RAG) 0.609 $\pm$ 0.018 0.661 $\pm$ 0.039 0.633 $\pm$ 0.024
AgentD (DeepSeek-Chat, w/ RAG) 0.685 $\pm$ 0.064 0.717 $\pm$ 0.042 0.700 $\pm$ 0.051
Vanilla DeepSeek-Chat (w/o RAG) 0.645 $\pm$ 0.021 0.698 $\pm$ 0.020 0.670 $\pm$ 0.012
To quantitatively assess these observed differences, we measured the semantic similarity between responses generated with and without RAG and human-annotated reference answers derived from the literature using BERTScore Zhang2019-nn. The reference answers were curated by the authors through careful reading of the relevant source papers qna_ex1, Leonetti2019. BERTScore quantifies contextual similarity by comparing the embeddings of model-generated and reference texts. We employed SciBERT Beltagy2019-eb as the embedding encoder because it is pre-trained on a large corpus of scientific publications from Semantic Scholar, capturing richer representations of biomedical and technical terminology than general-domain encoders such as RoBERTa Li2025-iy.
As shown in Table 2, AgentD consistently outperforms the corresponding vanilla language models across both GPT-4o and DeepSeek-Chat backends. For GPT-4o, the use of RAG increases the mean F1-score from 0.63 $\pm$ 0.01 to 0.70 $\pm$ 0.05, primarily driven by a substantial gain in Precision (0.72 vs. 0.61). This indicates that AgentD more effectively suppresses unsupported or generic statements during question answering. A similar trend is observed with DeepSeek-Chat: AgentD improves the F1-score from 0.67 $\pm$ 0.01 in the vanilla model to 0.70 $\pm$ 0.05, again with a noticeable increase in Precision (0.69 vs. 0.65). These results show that integrating RAG and structured reasoning enables AgentD to produce more faithful and literature-grounded responses.